# Variation of cones of divisors in a family of varieties – Fano type case

### v3

@@ 4.1. Deformation of Néron-Severi spaces
In the following, we show that if $X\to T$ is a family of Fano type varieties, then there exists a generic finite base change $T^{\prime}\to T$ and an open subset $U\subset T^{\prime}$, such that ${\mathcal{P}}ic_{X_{U}/U}\otimes\mathbb{R}$ is a constant sheaf. To be precise, this means that there exist natural isomorphisms
@@
First, recall the definition of the triviality of a map (see [Mat22, §11] for details). Consider a diagram of topological spaces:
${X}$${Y}$${Z.}$$\scriptstyle{f}$$\scriptstyle{\pi_{1}}$$\scriptstyle{\pi_{2}}$
@@
Suppose that $g:Y\to X$ is a birational morphism and $(X,\Delta)$ is a klt pair for some effective $\mathbb{R}$-divisor $\Delta$ on $X$. If $D$ is a $\mathbb{Q}$-Cartier divisor on $Y$ with $D\equiv 0/X$, then $g_{*}D$ is a $\mathbb{Q}$-Cartier divisor on $X$ such that $g^{*}(g_{*}D)=D$.
@@
Step 1. We prove the claim under the additional assumption that $X$ is smooth. We claim that there exists a generic finite base change $T^{\prime}\to T$ and an open subset $T_{0}\subset T^{\prime}$ such that
@@
for very general $t\in T_{0}$. This can be proved by the same argument as [KM92, Proposition 12.2.5]. Indeed, let $\mathcal{H}$ be the Hilbert scheme that parametrizes $(n-1)$-cycles of fibers of $f$, where $\dim(X/T)=n$. let $\mathcal{H}=\cup_{i\in J}\mathcal{H}_{i}$ be the decomposition into irreducible components. Then $J$ is at most a countable set and the natural map $p_{i}:\mathcal{H}_{i}\to T$ is projective. Let $I=\{i\in J\mid p_{i}(\mathcal{H}_{i})\neq T\}$. Then
@@
Shrinking $T$, we can assume that $X\to T$ is a locally trivial family in the analytic topology. As $X$ is smooth, shrinking $T$ further, we can assume that $f$ is a smooth morphism. Hence, $h^{1,1}(X_{t},\mathbb{Q})$ is invariant for $t\in T$. Fix some $t_{0}\in W$. By the choice of $W$, for each prime divisor $D$ on $X_{t_{0}}$, there exists some $i\in I$ and a Weil divisor $\mathcal{D}$ on $X_{\mathcal{H}_{i}}\coloneqq\mathcal{H}_{i}\times_{T}X$ such that $\mathcal{D}|_{X_{\mathcal{H}_{i},s_{0}}}=D$, where $s_{0}\in\mathcal{H}_{i}$ is a closed point such that $p_{i}(s_{0})=t_{0}$. Note that we have $X_{\mathcal{H}_{i},s_{0}}\simeq X_{t_{0}}$. Let $\mathcal{G}_{i}\subset\mathcal{H}_{i}$ be a closed subvariety such that $\mathcal{G}_{i}\to T$ is a generic finite morphism. Then we still have $\mathcal{D}|_{X_{\mathcal{G}_{i},s_{0}}}=D$, where $s_{0}\in\mathcal{G}_{i}$ is a closed point such that $p_{i}(s_{0})=t_{0}$. As $f$ is smooth, $f_{\mathcal{G}_{i}}$ is still a smooth morphism. Shrinking $\mathcal{G}_{i}$, we can assume that $\mathcal{G}_{i}$ is a smooth variety, and thus $X_{\mathcal{G}_{i}}$ is a smooth variety. Hence $\mathcal{D}|_{\mathcal{G}_{i}}$ is Cartier divisor on $X_{\mathcal{G}_{i}}$.
As $h^{1,1}(X_{t_{0}},\mathbb{Q})$ is of finite dimension, we can repeat the above process finitely many times, and obtain a generic finite morphism $q:T^{\prime}\to T$ and an open subset $T_{0}\subset T^{\prime}$ such that
@@
is surjective. Moreover, each Cartier divisor $B$ on $X_{T_{0},t_{0}}$ is a restriction of a cycle $\mathcal{B}$ on $X_{T_{0}}$ which is flat over $T_{0}$. Note that $X\to T$ is a locally trivial family in the analytic topology. Suppose that $t_{0}\in\mathbb{D}\subset T_{0}$ is a contractible analytic open set such that $X_{\mathbb{D}}$ is trivial. By the flatness of $\mathcal{B}$ over $T_{0}$, for each $t\in\mathbb{D}$, $[\mathcal{B}_{t}]$ corresponds to $[\mathcal{B}_{t_{0}}]$ under the identification $H^{2}(X_{t_{0}},\mathbb{Q})\simeq H^{2}(X_{t},\mathbb{Q})$. By the connectedness of $T_{0}$, we see that
@@
Let $h:Y\to X$ be a resolution which can be assumed to be a fiber-wise resolution after shrinking $T$. As $(X,\Delta)$ has klt singularities, after shrinking $T$, we can assume that $(X_{s},\Delta_{s})$ still has klt singularties for $s\in S$. In particular, $X_{s}$ has rational singularities (see [KM98, Theorem 5.22]). Thus, possibly shrinking $T$ further, $H^{i}(Y_{s},\mathcal{O}_{Y_{s}})=0,i=1,2$ for each $s\in S$. By Step 1, there exist a generic finite base change $q:T^{\prime}\to T$ and an open subset $T_{0}\subset T^{\prime}$, such that $N^{1}(Y_{T_{0}}/T_{0})\to N^{1}(Y_{T_{0},t})$ is surjective for $t\in q^{-1}(W)\cap T_{0}$, where $W\subset T$ consists of very general points. Moreover, each Cartier divisor $B$ on $Y_{T_{0},t_{0}}$ is a restriction of a cycle $\mathcal{B}$ on $Y_{T_{0}}$ which is flat over $T_{0}$.
Shrinking $T_{0}$, we can assume that $T_{0}\to q(T_{0})$ is an étale morphism. Hence there also exists a divisor $\Delta^{\prime}$ on $X_{T_{0}}$ such that $(X_{T_{0}},\Delta^{\prime})$ has klt singularities (see [KM98, Proposition 5.20]).
@@
By Lemma 4.2, we have $\mathcal{B}_{i}\sim_{\mathbb{Q}}0/X_{T_{0}}$. As $h_{T_{0}}$ is a birational morphism, $\mathcal{D}_{i}\coloneqq h_{T_{0},*}\mathcal{B}_{i}$ is a $\mathbb{Q}$-Cartier divisor and $\mathcal{B}_{i}=h_{T_{0}}^{*}\mathcal{D}_{i}$ by the negativity lemma. Therefore, we have
@@
Let $f:X\to T$ be a projective fibration with $(X,\Delta)$ a klt pair for some effective $\mathbb{R}$-divisor $\Delta$. Suppose that $S\subset T$ is a Zariski dense subset such that for any $s\in S$, the fiber $X_{s}$ is a rationally chain-connected variety. Then, up to a generic finite base change of $T$, there exists a non-empty open subset $T_{0}\subset T$, such that ${\mathcal{P}}ic_{X_{T_{0}}/T_{0}}\otimes\mathbb{R}$ is a constant sheaf. Moreover, for any open subset $U\subset T_{0}$, the natural restriction maps
@@
Shrinking $T$, we have $H^{1}(X_{s},\mathcal{O}_{X_{s}})=H^{2}(X_{s},\mathcal{O}_{X_{s}})=0$ for each $s\in S$ as $X_{s}$ is a rationally chain-connected variety and $(X_{s},\Delta_{s})$ has klt singularities. Hence, the claim follows from Theorem 1.3. The last claim follows as a Fano type variety is a rationally chain-connected variety. ∎
@@ 4.2. Deformation of nef cones
    for any $\mathcal{D}\in\overline{\rm{Eff}}(X_{V}/V)$, there exsits a $\mathcal{D}$-MMP$/V$ which terminates to a model such that the push-foreard of $\mathcal{D}$ is nef.
@@
$\mathcal{N}_{\mathcal{D}}\coloneqq\{t\in T\mid\mathcal{D}_{t}\text{~{} is nef~{}}\}.$
We claim that $\mathcal{N}$ is an open subset. By assumption, we can run some $\mathcal{D}$-MMP/$T$, $X\dasharrow X^{\prime}$, such that $\mathcal{D}^{\prime}$ is nef$/T$, where $\mathcal{D}^{\prime}$ is the strict transform of $\mathcal{D}$ on $X^{\prime}$. This MMP is invariant when restricting to each fiber $X_{t},t\in\mathcal{N}_{\mathcal{D}}$. As the image $Z\subset T$ of the contracting/flipping loci is a closed subset of $T$, and since $Z=T\setminus\mathcal{N}_{\mathcal{D}}$, it follows that $\mathcal{N}_{\mathcal{D}}$ is an open subset of $T$.
@@
are injective as $N^{1}(X_{T_{0}}/T_{0})\simeq N^{1}(X_{U}/U)\simeq N^{1}(X_{t})$. Thus, we have ${\rm{NE}}(X_{t})\subset{\rm{NE}}(X_{U}/U)$ under this identification. If the inclusion is strict, then there exists a divisor $\mathcal{D}$ on $X_{U}/U$ such that $\mathcal{D}$ is not nef$/U$ while $\mathcal{D}_{t}\in{\rm{Nef}}(X_{t})$. By ${\rm{Nef}}(X_{U}/U)\simeq{\rm{Nef}}(X_{t})$, we see that $\mathcal{D}$ is nef$/U$. This is a contradiction. The remaining isomorphisms can be shown similarly. ∎
@@
As $X_{t}$ is a Mori dream space, ${\rm{Nef}}(X_{t})$ is a rational polyhedral cone. As $X_{V}$ is a Mori dream space over $V$, for any pseudo-effective divisor $\mathcal{D}$ on $X_{V}$ over $V$, there exists a $\mathcal{D}$-MMP$/V$ that terminates to a model such that the push-forward of $\mathcal{D}$ is semi-ample. This verifies the conditions in Theorem 4.4, and thus the claim follows. ∎
@@
Let $f:X\to T$ be a projective fibration. Suppose that there exists a Zariski dense subset $S\subset T$ such that the fibers $X_{s}$ for $s\in S$ satisfy one of the cases in Theorem 1.2. Then, up to a generic finite base change of $T$, there exists a non-empty Zariski open subset $T_{0}\subset T$, such that for any open subset $U\subset T_{0}$, we have
@@
By Theorem 1.2, we can assume that $X$ is of Fano type over $T$ after shrinking $T$. Shrinking $T$ further, we can assume that $X_{t}$ is of Fano type for each $t\in T$. Note that a variety of Fano type (resp. over $T$) is a Mori dream space (resp. over $T$). Taking a generic finite base change of $T$ and shrinking $T$ again, we may assume that Corollary 4.5 (1) is satisfied by Theorem 1.3 (2). Therefore, the claim follows from Corollary 4.5. ∎

### v4

@@ 4.1. Deformation of Néron-Severi spaces
In the following, we show that if $X\to T$ is a family of Fano type varieties, then there exists a generically finite base change $T^{\prime}\to T$ and an open subset $U\subset T^{\prime}$, such that ${\mathcal{P}}ic_{X_{U}/U}\otimes\mathbb{R}$ is a constant sheaf. To be precise, this means that there exist natural isomorphisms
@@
First, recall the definition of the triviality of a map (see [Mat22, §11] for details). Consider a diagram of spaces and mappings in the analytic topology:
[…]
@@
Let $g:Y\to X$ be a projective morphism between normal varieties with $R^{1}g_{*}\mathcal{O}_{Y}=0$. If $D$ is a $\mathbb{Q}$-Cartier divisor on $Y$ with $D\equiv 0/X$, then $D\sim_{\mathbb{Q}}0/X$.
In particular, suppose that $g$ is a birational morphism and that $(X,\Delta)$ is a klt pair for some effective $\mathbb{R}$-divisor $\Delta$ on $X$. If $D$ is a $\mathbb{Q}$-Cartier divisor on $Y$ with $D\equiv 0/X$, then $g_{*}D$ is a $\mathbb{Q}$-Cartier divisor on $X$ such that $g^{*}(g_{*}D)=D$.
@@
The argument below closely follows the proof of [KM92, Proposition (12.1.4)] and strengthens it. For the reader’s convenience, we include the full proof.
Let $x\in X$ be a point. Then there exists a triangulation of the topological space $Y$ such that the analytic subset $g^{-1}(x)$ appears as the support of a sub-complex (see [BHPV04, I, Theorem (8.8)]). Let $x\in U$ be a sufficiently small Stein neighborhood. Then $g^{-1}(U)$ deformation retracts onto the sub-complex $g^{-1}(x)$ (see [Hat00, Appendix, Proposition A.5]). In particular, we have
$H^{i}(g^{-1}(U),\mathbb{Z})=H^{i}(g^{-1}(x),\mathbb{Z}),i\in\mathbb{Z}_{\geq 0}.$
Let $(-)^{\rm an}$ denote the analytification functor. As $R^{1}g^{\rm an}_{*}\mathcal{O}^{\rm an}_{Y}=(R^{1}g_{*}\mathcal{O}_{Y})^{\rm an}=0$ (see [SGA1, Exposé XII, Théorème 4.2]) and $H^{1}(U,\mathcal{O}^{\rm an}_{X})=0$, we have
$H^{1}(g^{-1}(U),\mathcal{O}^{\rm an}_{Y})=0.$
By the exponential sequence, we have
$H^{1}(g^{-1}(U),\mathcal{O}^{\rm an}_{Y})\to H^{1}(g^{-1}(U),\mathcal{O}_{Y}^{\rm an,*})\to H^{2}(g^{-1}(U),\mathbb{Z}),$
and thus induces the natural injective group homomorphism
${{\rm{Pic}}}(g^{-1}(U))\hookrightarrow H^{2}(g^{-1}(U),\mathbb{Z})=H^{2}(g^{-1}(x),\mathbb{Z}).$
If $D\equiv 0/X$, then $D|_{g^{-1}(x)}\equiv 0$ and thus $D|_{g^{-1}(U)}$ is a torsion element of $H^{2}(g^{-1}(x),\mathbb{Z})$. If $d$ is the order of the torsion group of $H^{2}(g^{-1}(x),\mathbb{Z})$, then $dD|_{g^{-1}(U)}=0\in{\rm{Pic}}(g^{-1}(U))$. Note that there are only finitely many topological types of the fibers of $g$ (for example, using Lemma 4.1). We can take $d\in\mathbb{N}$ sufficiently divisible, such that for each $x^{\prime}\in X$, there exists an analytic neighborhood $x^{\prime}\in U^{\prime}$ such that $dD|_{g^{-1}{(U^{\prime})}}\sim 0$ on $g^{-1}{(U^{\prime})}$. As $g^{\rm an}_{*}\mathcal{O}^{\rm an}_{Y}(dD)\simeq(g_{*}\mathcal{O}_{Y}(dD))^{\rm an}$ (see [SGA1, Exposé XII, Théorème 4.2]), we have that $g_{*}\mathcal{O}_{Y}(dD)$ is a line bundle on $X$. By the natural sheaf morphism, $g^{*}g_{*}(\mathcal{O}_{Y}(dD))\to\mathcal{O}_{Y}(dD)$, we have the natural sheaf morphism (in the Zariski topology)
$\theta:\mathcal{O}_{Y}\to\mathcal{O}_{Y}(dD)\otimes(g^{*}g_{*}(\mathcal{O}_{Y}(dD)))^{\vee}.$
On $g^{-1}(U^{\prime})$, as $g^{\rm an,*}g^{\rm an}_{*}(\mathcal{O}^{\rm an}_{Y}(dD))\to\mathcal{O}^{\rm an}_{Y}(dD)$ is surjective and $\mathcal{O}^{\rm an}_{Y}$ is torsion free, we see that $\theta^{\rm an}|_{g^{-1}(U^{\prime})}$ is an isomorphism. Thus $\theta^{\rm an}$ is an isomorphism. This implies that $\theta$ is an isomorphism (for example see [SGA1, Exposé XII, Théorème et définition 1.1]). This shows $D\sim_{\mathbb{Q}}0/X$.
Next, assume that $g$ is birational and there exists a divisor $\Delta$ such that $(X,\Delta)$ has klt singularities. Replacing $Y$ by a resolution, we can assume that $Y$ is smooth. As klt singularities are rational singularities, we have $R^{1}g_{*}\mathcal{O}_{Y}=0$. Hence, we have $D\sim_{\mathbb{Q}}0/X$ by the previous discussion. This implies that $g_{*}D$ is a $\mathbb{Q}$-Cartier divisor on $X$ and $D=g^{*}(g_{*}D)$ by the negativity lemma. This completes the proof.
Finally, we include an alternative proof of this special case below, which is more algebraic in nature.
@@
Remark 4.3.
Lemma 4.2 does not hold for an arbitrary birational contraction $g$. Assume that $X$ is a normal projective surface which is not $\mathbb{Q}$-factorial. Let $H$ be a Weil divisor on $X$ which is not $\mathbb{Q}$-Cartier. For a resolution $g:Y\to X$, one can define the Mumford pull-back of $H$, denoted by $g^{*}H$. Then we have $g^{*}H\equiv 0/X$ by construction. However, we do not have $g^{*}H\sim_{\mathbb{Q}}0/X$, since this would imply that $H$ is $\mathbb{Q}$-Cartier. Lemma 4.2 is consistent with the fact that for projective surfaces, rational singularities are $\mathbb{Q}$-factorial (see [Bad01, Theorem 4.6]).
@@
Step 1. We prove the claim under the additional assumption that $X$ is smooth. We claim that there exists a generically finite base change $T^{\prime}\to T$ and an open subset $T_{0}\subset T^{\prime}$ such that
@@
for very general $t\in T_{0}$. This can be proved by the same argument as [KM92, Proposition 12.2.5]. Indeed, let $\mathcal{H}$ be the Hilbert scheme that parametrizes (n1)(n-1)-cycles of fibers of $f$, where $\dim(X/T)=n$. let $\mathcal{H}=\cup_{i\in J}\mathcal{H}_{i}$ be the decomposition into irreducible components. Then $J$ is at most a countable set and the natural map $p_{i}:\mathcal{H}_{i}\to T$ is projective. Let $I=\{i\in J\mid p_{i}(\mathcal{H}_{i})\neq T\}$. Then
@@
Shrinking $T$, we can assume that $X\to T$ is a locally trivial family in the analytic topology. As $X$ is smooth, shrinking $T$ further, we can assume that $f$ is a smooth morphism. Hence, $h^{1,1}(X_{t},\mathbb{Q})$ is invariant for $t\in T$. Fix some $t_{0}\in W$. By the choice of $W$, for each prime divisor $D$ on $X_{t_{0}}$, there exists some $i\in I$ and a Weil divisor $\mathcal{D}$ on $X_{\mathcal{H}_{i}}\coloneqq\mathcal{H}_{i}\times_{T}X$ such that $\mathcal{D}|_{X_{\mathcal{H}_{i},s_{0}}}=D$, where $s_{0}\in\mathcal{H}_{i}$ is a closed point such that $p_{i}(s_{0})=t_{0}$. Note that we have $X_{\mathcal{H}_{i},s_{0}}\simeq X_{t_{0}}$. Let $\mathcal{G}_{i}\subset\mathcal{H}_{i}$ be a closed subvariety such that $\mathcal{G}_{i}\to T$ is a generically finite morphism. Then we still have $\mathcal{D}|_{X_{\mathcal{G}_{i},s_{0}}}=D$, where $s_{0}\in\mathcal{G}_{i}$ is a closed point such that $p_{i}(s_{0})=t_{0}$. As $f$ is smooth, $f_{\mathcal{G}_{i}}$ is still a smooth morphism. Shrinking $\mathcal{G}_{i}$, we can assume that $\mathcal{G}_{i}$ is a smooth variety, and thus $X_{\mathcal{G}_{i}}$ is a smooth variety. Hence $\mathcal{D}|_{\mathcal{G}_{i}}$ is Cartier divisor on $X_{\mathcal{G}_{i}}$.
As $h^{1,1}(X_{t_{0}},\mathbb{Q})$ is of finite dimension, we can repeat the above process finitely many times, and obtain a generically finite morphism $q:T^{\prime}\to T$ and an open subset $T_{0}\subset T^{\prime}$ such that
@@
is surjective. Moreover, each Cartier divisor $B$ on $X_{T_{0},t_{0}}$ is a restriction of a cycle $\mathcal{B}$ on $X_{T_{0}}$ which is flat over $T_{0}$. Note that $X\to T$ is a locally trivial family in the analytic topology. Suppose that $t_{0}\in\mathbb{D}\subset T_{0}$ is a contractible analytic open set such that $X_{\mathbb{D}}$ is trivial. By the flatness of $\mathcal{B}$ over $T_{0}$, for each $\in\mathbb{D}, [t][\mathcal{B}_{t}$] corresponds to $\mathcal{B}_{t_{0}}]$ under the identification $H^{2}(X_{t_{0}},\mathbb{Q})\simeq H^{2}(X_{t},\mathbb{Q})$. By the connectedness of $T_{0}$, we see that
@@
Let $h:Y\to X$ be a resolution which can be assumed to be a fiber-wise resolution after shrinking $T$. Note that for the dense subset $S$, each point $s\in S$ satisfies that the fiber $Y_{s}$ has rational singularities and $H^{i}(Y_{s},\mathcal{O}_{Y_{s}})=0,i=1,2$. By Step 1, there exist a generically finite base change $q:T^{\prime}\to T$ and an open subset $T_{0}\subset T^{\prime}$, such that $N^{1}(Y_{T_{0}}/T_{0})\to N^{1}(Y_{T_{0},t})$ is surjective for $t\in q^{-1}(W)\cap T_{0}$, where $W\subset T$ consists of very general points. Moreover, each Cartier divisor $B$ on $Y_{T_{0},t_{0}}$ is a restriction of a cycle $\mathcal{B}$ on $Y_{T_{0}}$ which is flat over $T_{0}$.
@@
As general fibers of $f_{T_{0}}$ have rational singularities, after shrinking $T_{0}$, we can assume that $X_{T_{0}}$ still have rational singularities (see [Elk78]). By Lemma 4.2, we have $\mathcal{B}_{i}\sim_{\mathbb{Q}}0/X_{T_{0}}$. As $h_{T_{0}}$ is a birational morphism, $\mathcal{D}_{i}\coloneqq h_{T_{0},*}\mathcal{B}_{i}$ is a $\mathbb{Q}$-Cartier divisor and $\mathcal{B}_{i}=h_{T_{0}}^{*}\mathcal{D}_{i}$ by the negativity lemma. Therefore, we have
@@
Let $f:X\to T$ be a projective fibration. Suppose that $S\subset T$ is a Zariski dense subset such that for any $s\in S$, the fiber $X_{s}$ is a rationally chain-connected variety with rational singularities. Then, up to a generically finite base change of $T$, there exists a non-empty open subset $T_{0}\subset T$, such that ${\mathcal{P}}ic_{X_{T_{0}}/T_{0}}\otimes\mathbb{R}$ is a constant sheaf. Moreover, for any open subset $U\subset T_{0}$, the natural restriction maps
@@
Shrinking $T$, we have $H^{1}(X_{s},\mathcal{O}_{X_{s}})=H^{2}(X_{s},\mathcal{O}_{X_{s}})=0$ for each $s\in S$ as $X_{s}$ is a rationally chain-connected variety with rational singularities. Hence, the claim follows from Theorem 1.3. Besides, a Fano type variety is naturally a rationally chain-connected variety with rational singularities. ∎
@@ 4.2. Deformation of nef cones
    for any $\mathcal{D}\in\overline{\rm{Eff}}(X_{V}/V)$, there exsits a $\mathcal{D}$-MMP/V/V which terminates to a model such that the push-foreard of $\mathcal{D}$ is nef.
@@
[…]
We claim that $\mathcal{N}$ is an open subset. By assumption, we can run some $\mathcal{D}$-MMP/$T$, $X\dasharrow X^{\prime}$, such that $\mathcal{D}^{\prime}$ is nef/T/T, where $\mathcal{D}^{\prime}$ is the strict transform of $\mathcal{D}$ on $X^{\prime}$. This MMP is invariant when restricting to each fiber $X_{t},t\in\mathcal{N}_{\mathcal{D}}$. As the image $Z\subset T$ of the contracting/flipping loci is a closed subset of $T$, and since $Z=T\setminus\mathcal{N}_{\mathcal{D}}$, it follows that $\mathcal{N}_{\mathcal{D}}$ is an open subset of $T$.
@@
are injective as $N^{1}(X_{T_{0}}/T_{0})\simeq N^{1}(X_{U}/U)\simeq N^{1}(X_{t})$. Thus, we have ${\rm{NE}}(X_{t})\subset{\rm{NE}}(X_{U}/U)$ under this identification. If the inclusion is strict, then there exists a divisor $\mathcal{D}$ on $X_{U}/U$ such that $\mathcal{D}$ is not nef/U/U while $\mathcal{D}_{t}\in{\rm{Nef}}(X_{t})$. By ${\rm{Nef}}(X_{U}/U)\simeq{\rm{Nef}}(X_{t})$, we see that $\mathcal{D}$ is nef/U/U. This is a contradiction. The remaining isomorphisms can be shown similarly. ∎
@@
As $X_{t}$ is a Mori dream space, ${\rm{Nef}}(X_{t})$ is a rational polyhedral cone. As $X_{V}$ is a Mori dream space over $V$, for any pseudo-effective divisor $\mathcal{D}$ on $X_{V}$ over $V$, there exists a $\mathcal{D}$-MMP/V/V that terminates to a model such that the push-forward of $\mathcal{D}$ is semi-ample. This verifies the conditions in Theorem 4.5, and thus the claim follows. ∎
@@
Let $f:X\to T$ be a projective fibration. Suppose that there exists a Zariski dense subset $S\subset T$ such that the fibers $X_{s}$ for $s\in S$ satisfy one of the cases in Theorem 1.2. Then, up to a generically finite base change of $T$, there exists a non-empty Zariski open subset $T_{0}\subset T$, such that for any open subset $U\subset T_{0}$, we have
@@
By Theorem 1.2, we can assume that $X$ is of Fano type over $T$ after shrinking $T$. Shrinking $T$ further, we can assume that $X_{t}$ is of Fano type for each $t\in T$. Note that a variety of Fano type (resp. over $T$) is a Mori dream space (resp. over $T$). Taking a generically finite base change of $T$ and shrinking $T$ again, we may assume that Corollary 4.6 (1) is satisfied by Theorem 1.3 (2). Therefore, the claim follows from Corollary 4.6. ∎